Two spherical objects have masses of 8000 kg and 1500 kg.  Their centers are separated by a distance of 1.5 m.  Find the gravitational attraction between them.


1
Expert's answer
2021-03-08T07:29:14-0500
"F=\dfrac{Gm_1m_2}{r^2},""F=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot8000\ kg\cdot1500\ kg}{(1.5\ m)^2}=3.56\cdot10^{-4}\ N."